# LoRA-FAIR: Federated LoRA Fine-Tuning with Aggregation and Initialization Refinement

### v2

@@ 4 LoRA-FAIR: Simple but Effective Solution
$\arg\min\limits_{\mathbf{\Delta B}}\underbrace{\mathcal{S}\left(\mathbf{\Delta W},(\mathbf{\bar{B}}+\mathbf{\Delta B})\mathbf{\bar{A}}\right)}_{\text{correction}}+\underbrace{\lambda||\mathbf{\Delta B}||}_{\text{regularization}},$ (8)
@@ 5.2 Ablation Studies
Impact of Regularization Weight $\lambda$. In our proposed method, we optimize the objective in Eq. 8 to address both server aggregation bias and client initialization lag. Notably, we include a regularization term $\lambda||\mathbf{\Delta B}||$ to balance the similarity measure with the correction term. Here, we conduct experiments to investigate the impact of the regularization weight $\lambda$ on model performance. As shown in Fig. 6, varying $\lambda$ affects the performance of LoRA-FAIR, highlighting the importance of this parameter. Specifically, when $\lambda=0$, LoRA-FAIR achieves its lowest performance.
Regularization Term $\lambda||\mathbf{\Delta B}||$ $\mathcal{S}(\mathbf{\bar{B}},\mathbf{\bar{B}}+\mathbf{\Delta B})$ $\mathcal{S}\left(\mathbf{\Delta W},(\mathbf{\bar{B}}+\mathbf{\Delta B})\mathbf{\bar{A}}\right)$ Average Accuracy
@@ 6 Related Work
Federated Learning. FedAvg [24], the foundational work in FL, demonstrates the advantages of this approach in terms of privacy and communication efficiency by aggregating local model parameters to train a shared global model. Numerous FL studies [24, 39, 42, 38, 37, 3, 27] have addressed various challenges within FL settings. For example, several works explore the impact of different initialization strategies on model performance. [32] shows that initializing with pre-trained weights can enhance the stability of FedAvg’s global aggregation, while [34] confirms the effectiveness of using a pre-trained model as an initial starting point. However, these methods primarily focus on smaller models and do not extend to foundation models or incorporate parameter-efficient fine-tuning; instead, they adhere to conventional FL training practices.
Federated Fine-Tuning. Several studies [7, 17, 41, 31, 40, 2] have explored federated fine-tuning approaches. For example, Kuang et al. [17] proposes federated fine-tuning with all parameters updated, while Sun et al. [30] introduces federated fine-tuning with PEFT using prefix-tuning. A closely related area to our work involves federated fine-tuning using LoRA. Zhang et al. [44] is the first study to apply LoRA in a federated context; however, this method overlooks potential server aggregation bias. Several subsequent works have been proposed: FFA-LoRA [31] freezes the non-zero initialized low-rank matrices and updates only the zero-initialized matrices, FlexLoRA [2] uses SVD to redistribute weights, and FLoRA [40] stacks local LoRA modules and transmits them to each client. However, these methods do not address client initialization lag. A more detailed discussion of related federated LoRA fine-tuning works can be found in Appendix Sec. 9.

### v3

@@ 4 LoRA-FAIR: Simple but Effective Solution
[…]
@@ 5.2 Ablation Studies
[…]
@@ 6 Related Work
Federated Learning. FedAvg [26], the foundational work in FL, demonstrates the advantages of this approach in terms of privacy and communication efficiency by aggregating local model parameters to train a shared global model. Numerous FL studies [26, 41, 44, 40, 39, 3, 29, 23, 24] have addressed various challenges within FL settings. For example, several works explore the impact of different initialization strategies on model performance. [34] shows that initializing with pre-trained weights can enhance the stability of FedAvg’s global aggregation, while [36] confirms the effectiveness of using a pre-trained model as an initial starting point. However, these methods primarily focus on smaller models and do not extend to foundation models or incorporate parameter-efficient fine-tuning; instead, they adhere to conventional FL training practices.
Federated Fine-Tuning. Several studies [7, 17, 43, 33, 42, 2] have explored federated fine-tuning approaches. For example, Kuang et al. [17] proposes federated fine-tuning with all parameters updated, while Sun et al. [32] introduces federated fine-tuning with PEFT using prefix-tuning. A closely related area to our work involves federated fine-tuning using LoRA. Zhang et al. [46, 47] apply LoRA in a federated context; however, these methods overlook potential server aggregation bias. Several subsequent works have been proposed: FFA-LoRA [33] freezes the non-zero initialized low-rank matrices and updates only the zero-initialized matrices, FlexLoRA [2] uses SVD to redistribute weights, and FLoRA [42] stacks local LoRA modules and transmits them to each client. However, these methods do not address client initialization lag. A more detailed discussion of related federated LoRA works can be found in Appendix Sec. 10.
@@ 7 Conclusion
8 Acknowledgments
The work of J. Bian, L. Wang and J. Xu is partially supported by NSF under grants 2505381 and 2515982. The work of L. Zhang is partially supported by NSF under grant 2348279.
@@ 10 Prior Works
11 Theorem
Theorem 11.1.
For analytical tractability, we consider the case where the similarity metric S is based on the Frobenius norm. The residual correction term $\Delta\mathbf{B}$ obtained by minimizing Equation (8) guarantees that $(\bar{\mathbf{B}}+\Delta\mathbf{B})\bar{\mathbf{A}}$ approaches the ideal global update $\Delta\mathbf{W}$ with the following approximation guarantee:
$\|(\bar{\mathbf{B}}+\Delta\mathbf{B}^{*})\bar{\mathbf{A}}-\Delta\mathbf{W}\|_{F}^{2}$
$\leq$ $\|\Delta\mathbf{W}-\bar{\mathbf{B}}\bar{\mathbf{A}}\|_{F}^{2}\cdot\left(1-\frac{\sigma_{\min}^{2}(\bar{\mathbf{A}})}{\sigma_{\min}^{2}(\bar{\mathbf{A}})+\lambda}\right)^{2},$ (9)
here $\sigma_{\min}(\bar{\mathbf{A}})$ is the smallest non-zero singular value of $\bar{\mathbf{A}}$.
Proof.
Let’s denote $\mathbf{E}=\Delta\mathbf{W}-\bar{\mathbf{B}}\bar{\mathbf{A}}$ as the initial aggregation error. Our objective function becomes:
$J(\Delta\mathbf{B})=\|\Delta\mathbf{B}\bar{\mathbf{A}}-\mathbf{E}\|_{F}^{2}+\lambda\|\Delta\mathbf{B}\|_{F}^{2}$ (10)
To find the critical points of $J(\Delta\mathbf{B})$, we take the derivative with respect to $\Delta\mathbf{B}$ and set it equal to zero:
$\nabla_{\Delta\mathbf{B}}J(\Delta\mathbf{B})=2(\Delta\mathbf{B}\bar{\mathbf{A}}-\mathbf{E})\bar{\mathbf{A}}^{T}+2\lambda\Delta\mathbf{B}=0$ (11)
Solving for the optimal $\Delta\mathbf{B}^{*}$:
$\Delta\mathbf{B}^{*}=\mathbf{E}\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1},$ (12)
where $\mathbf{I}$ is the identity matrix of appropriate dimensions. Substituting back the definition of $\mathbf{E}$:
$\Delta\mathbf{B}^{*}=(\Delta\mathbf{W}-\bar{\mathbf{B}}\bar{\mathbf{A}})\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}$ (13)
The residual error after applying the correction is:
$\mathbf{E}_{residual}$ $=-\mathbf{E}+\Delta\mathbf{B}^{*}\bar{\mathbf{A}}$
$=-\mathbf{E}+\mathbf{E}\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}\bar{\mathbf{A}}$
$=\mathbf{E}(-\mathbf{I}+\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}\bar{\mathbf{A}})$ (14)
Let’s define the matrix $\mathbf{M}=-\mathbf{I}+\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}\bar{\mathbf{A}}$. For any matrix $\bar{\mathbf{A}}$, the eigenvalues of $\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}\bar{\mathbf{A}}$ can be bounded using the properties of matrix norms and the Sherman-Morrison-Woodbury formula:
$\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}\bar{\mathbf{A}}=\bar{\mathbf{A}}^{T}\bar{\mathbf{A}}(\bar{\mathbf{A}}^{T}\bar{\mathbf{A}}+\lambda\mathbf{I})^{-1}$ (15)
The eigenvalues of this matrix are of the form $\frac{\mu_{i}}{\mu_{i}+\lambda}$, where $\mu_{i}$ are the eigenvalues of $\bar{\mathbf{A}}^{T}\bar{\mathbf{A}}$. Since the eigenvalues of $\bar{\mathbf{A}}^{T}\bar{\mathbf{A}}$ are the squares of the singular values of $\bar{\mathbf{A}}$, i.e., $\mu_{i}=\sigma_{i}^{2}$, the eigenvalues of $\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}\bar{\mathbf{A}}$ are $\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}+\lambda}$. Therefore, the eigenvalues of $\mathbf{M}=-\mathbf{I}+\bar{\mathbf{A}}^{T}(\bar{\mathbf{A}}\bar{\mathbf{A}}^{T}+\lambda\mathbf{I})^{-1}\bar{\mathbf{A}}$ are $1+\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}+\lambda}=-\frac{\lambda}{\sigma_{i}^{2}+\lambda}$.
The spectral norm of $\mathbf{M}$ is the maximum absolute eigenvalue:
$\|\mathbf{M}\|_{2}=\max_{i}\left|\frac{-\lambda}{\sigma_{i}^{2}+\lambda}\right|=\frac{\lambda}{\sigma_{\min}^{2}+\lambda}$ (16)
Using the property that for any matrices $\mathbf{P}$ and $\mathbf{Q}$, $\|\mathbf{P}\mathbf{Q}\|_{F}\leq\|\mathbf{P}\|_{F}\|\mathbf{Q}\|_{2}$, we have:
$\|\mathbf{E}_{residual}\|_{F}$ $=\|\mathbf{E}\mathbf{M}\|_{F}$ (17)
$\leq\|\mathbf{E}\|_{F}\|\mathbf{M}\|_{2}$ (18)
$=\|\mathbf{E}\|_{F}\cdot\frac{\lambda}{\sigma_{\min}^{2}+\lambda}$ (19)
Since $\frac{\lambda}{\sigma_{\min}^{2}+\lambda}=1-\frac{\sigma_{\min}^{2}}{\sigma_{\min}^{2}+\lambda}$, we have:
$\|\mathbf{E}_{residual}\|_{F}\leq\|\mathbf{E}\|_{F}\cdot\left(1-\frac{\sigma_{\min}^{2}}{\sigma_{\min}^{2}+\lambda}\right)$ (20)
Squaring both sides and substituting $\mathbf{E}=\Delta\mathbf{W}-\bar{\mathbf{B}}\bar{\mathbf{A}}$, we get our final bound.
Corollary 11.2.
When $\lambda=0$ and $\bar{\mathbf{A}}$ has full row rank, there exists an exact solution where:
$(\bar{\mathbf{B}}+\Delta\mathbf{B}^{*})\bar{\mathbf{A}}=\Delta\mathbf{W}$
As the regularization parameter $\lambda$ increases, the solution balances two objectives:
  1. 1.
    Minimizing the approximation error to the ideal update $\Delta\mathbf{W}$
  2. 2.
    Preventing large deviations from the averaged LoRA module $\bar{\mathbf{B}}$
Theorem 11.3 (Convergence of Federated LoRA Fine-Tuning).
Under standard FL assumptions (L-smooth loss, bounded gradients $G$, $E$ local epochs), and assuming the global learning rate $\eta$, the convergence of federated LoRA fine-tuning after $T$ rounds is:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(W^{t})\|^{2}]\leq\frac{4[\mathcal{L}(W^{0})-\mathcal{L}(W^{*})]}{\eta T}$
$+4\eta^{2}E^{2}G^{2}\left(\frac{L^{2}}{2}+1\right)+8\frac{1}{T}\sum_{t=0}^{T-1}\|\Delta W^{t}-\bar{B}^{t}\bar{A}^{t}\|_{F}^{2}\cdot\gamma$
where $\gamma$ characterizes the aggregation method: 1. For LoRA-FAIR: $\gamma=\left(1-\frac{\sigma^{2}_{\min}(\bar{A}^{t})}{\sigma^{2}_{\min}(\bar{A}^{t})+\lambda}\right)^{2}<1$. 2. For FedIT (standard aggregation): $\gamma=1$
Proof.
At round $t$, the global model update is:
$W^{t+1}=W^{t}-\eta\cdot\Delta W^{\prime t}$ (21)
where $\Delta W^{\prime t}=(\bar{B}^{t}+\Delta B^{*t})\bar{A}^{t}$ for LoRA-FAIR and $\Delta W^{\prime t}=\bar{B}^{t}\bar{A}^{t}$ for FedIT. The ideal global update is:
$\Delta W^{t}=\sum_{k=1}^{K}p_{k}B_{k}^{t}A_{k}^{t}$ (22)
Define the aggregation error:
$E_{agg}^{t}=\Delta W^{t}-\Delta W^{\prime t}$ (23)
From Theorem A.1, for LoRA-FAIR:
$\|E_{agg}^{t}\|_{F}^{2}\leq\|\Delta W^{t}-\bar{B}^{t}\bar{A}^{t}\|_{F}^{2}\cdot\gamma$ (24)
Using L-smoothness:
$\mathcal{L}(W^{t+1})$ $\leq\mathcal{L}(W^{t})+\langle\nabla\mathcal{L}(W^{t}),W^{t+1}-W^{t}\rangle$
$\quad+\frac{L}{2}\|W^{t+1}-W^{t}\|^{2}$ (25)
$=\mathcal{L}(W^{t})-\eta\langle\nabla\mathcal{L}(W^{t}),\Delta W^{\prime t}\rangle$
$\quad+\frac{L\eta^{2}}{2}\|\Delta W^{\prime t}\|^{2}$ (26)
Decomposing $\Delta W^{\prime t}=\Delta W^{t}-E_{agg}^{t}$:
$\langle\nabla\mathcal{L}(W^{t}),\Delta W^{\prime t}\rangle$ $=\langle\nabla\mathcal{L}(W^{t}),\Delta W^{t}\rangle$
$\quad-\langle\nabla\mathcal{L}(W^{t}),E_{agg}^{t}\rangle$ (27)
Under bounded gradients:
$\langle\nabla\mathcal{L}(W^{t}),\Delta W^{t}\rangle\geq\frac{1}{2}\|\nabla\mathcal{L}(W^{t})\|^{2}-\frac{\eta^{2}E^{2}L^{2}G^{2}}{2}$ (28)
For the error term:
$\langle\nabla\mathcal{L}(W^{t}),-E_{agg}^{t}\rangle\leq\frac{1}{4}\|\nabla\mathcal{L}(W^{t})\|^{2}+\|E_{agg}^{t}\|^{2}$ (29)
Combining bounds:
$\mathcal{L}(W^{t+1})$ $\leq\mathcal{L}(W^{t})-\frac{\eta}{4}\|\nabla\mathcal{L}(W^{t})\|^{2}+\frac{\eta^{3}E^{2}L^{2}G^{2}}{2}$
$\quad+\eta\|E_{agg}^{t}\|^{2}+\frac{L\eta^{2}}{2}\|\Delta W^{\prime t}\|^{2}$ (30)
Using $\|\Delta W^{\prime t}\|^{2}\leq 2(\|\Delta W^{t}\|^{2}+\|E_{agg}^{t}\|^{2})$ and the bound on $E_{agg}^{t}$:
$\mathcal{L}(W^{t+1})$ $\leq\mathcal{L}(W^{t})-\frac{\eta}{4}\|\nabla\mathcal{L}(W^{t})\|^{2}+\frac{\eta^{3}E^{2}L^{2}G^{2}}{2}$
$\quad+\eta\|E_{agg}^{t}\|^{2}+\frac{L\eta^{2}}{2}(2(\|\Delta W^{t}\|^{2}+\|E_{agg}^{t}\|^{2}))$
$\leq\mathcal{L}(W^{t})-\frac{\eta}{4}\|\nabla\mathcal{L}(W^{t})\|^{2}+\frac{\eta^{3}E^{2}L^{2}G^{2}}{2}$
$\quad+\eta\|E_{agg}^{t}\|^{2}+{\eta}(\eta^{2}E^{2}G^{2}+\|E_{agg}^{t}\|^{2})$
$\leq\mathcal{L}(W^{t})-\frac{\eta}{4}\|\nabla\mathcal{L}(W^{t})\|^{2}+\frac{\eta^{3}E^{2}L^{2}G^{2}}{2}$
$\quad+2\eta\|E_{agg}^{t}\|^{2}+\eta^{3}E^{2}G^{2}$ (31)
where $\eta<\frac{1}{L}$. Then we have:
$\mathcal{L}(W^{t+1})$ $\leq\mathcal{L}(W^{t})-\frac{\eta}{4}\|\nabla\mathcal{L}(W^{t})\|^{2}+\frac{\eta^{3}E^{2}L^{2}G^{2}}{2}$
$\quad+2\eta\|\Delta W^{t}-\bar{B}^{t}\bar{A}^{t}\|_{F}^{2}\cdot\gamma+\eta^{3}E^{2}G^{2}$ (32)
Rearranging and taking expectation:
$\frac{\eta}{4}\mathbb{E}[\|\nabla\mathcal{L}(W^{t})\|^{2}]$ $\leq\mathbb{E}[\mathcal{L}(W^{t})]-\mathbb{E}[\mathcal{L}(W^{t+1})]$
$\quad+\frac{\eta^{3}E^{2}L^{2}G^{2}}{2}+\eta^{3}E^{2}G^{2}$
$\quad+2\eta\|\Delta W^{t}-\bar{B}^{t}\bar{A}^{t}\|_{F}^{2}\cdot\gamma$ (33)
Then by summing over t:
$\frac{\eta}{4}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(W^{t})\|^{2}]$ $\leq\mathcal{L}(W^{0})-\mathbb{E}[\mathcal{L}(W^{T})]$
$\quad+T\eta^{3}E^{2}G^{2}(\frac{L^{2}}{2}+1)$
$\quad+2\eta\sum_{t=0}^{T-1}\|\Delta W^{t}-\bar{B}^{t}\bar{A}^{t}\|_{F}^{2}\cdot\gamma$ (34)
dividing both sides by $\frac{\eta}{4}$ and T, we have:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(W^{t})\|^{2}]$ $\leq\frac{4}{T\eta}[\mathcal{L}(W^{0})-\mathcal{L}(W^{*})]$
$\quad+4\eta^{2}E^{2}G^{2}(\frac{L^{2}}{2}+1)$
$\quad+8\frac{1}{T}\sum_{t=0}^{T-1}\|\Delta W^{t}-\bar{B}^{t}\bar{A}^{t}\|_{F}^{2}\cdot\gamma$ (35)
With $\eta=\frac{1}{\sqrt{T}}$, we finally get:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(W^{t})\|^{2}]$ $\leq\frac{4[\mathcal{L}(W^{0})-\mathcal{L}(W^{*})]}{\sqrt{T}}$
$\quad+4\eta^{2}E^{2}G^{2}(\frac{L^{2}}{2}+1)$
$\quad+8\frac{1}{T}\sum_{t=0}^{T-1}\|\Delta W^{t}-\bar{B}^{t}\bar{A}^{t}\|_{F}^{2}\cdot\gamma$ (36)
As data becomes more non-IID, $\|\Delta\mathbf{W}-\bar{\mathbf{B}}\bar{\mathbf{A}}\|_{F}$ increases, leading to a larger error term in standard methods. LoRA-FAIR reduces this impact by the factor $\gamma<1$, providing tighter convergence bounds especially in highly non-IID settings.